OpenStudy (anonymous):

PLEASE HELP! Where does the graph of y = 2x2 − 5x + 3 cross the x-axis

OpenStudy (jkristia):

you have a quadratic equation and the easiest way to solve this is using the quadratic formula

OpenStudy (anonymous):

Okay. That didn't help. but i have to get off now.

OpenStudy (jkristia):

a quadratic equation is of the form \[ax^2 + bx + c = 0\] and x can be found using the quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a} \] does that help ?

OpenStudy (jkristia):

for more information take a look at wikipedia http://en.wikipedia.org/wiki/Quadratic_equation